# Quadratic-time computations for pseudo-Anosov mapping classes

### v1

@@ 1.2. The algorithm
There are many examples of ledgers already in the literature. In his thesis [26], Penner gave a ledger in terms of Dehn–Thurston train track coordinates (explicit matrices are not given, but they can be derived from the computations there). Penner’s formulas contain minor mistakes, corrected in his subsequent paper [23]
@@
The complexity of the sink package algorithm is quadratic in the bit length of the entries of $A$. Indeed, the Zassenhaus algorithm converts the problem of computing a basis for an intersection into a row reduction problem, and the complexity of row reduction for integer matrices is known to be quadratic in the bit lengths of the entries; see, e.g. [31, Lecture X, Theorem 5].
@@
We are finally ready to state our corollary of Theorem 1.1.
@@ 1.3. New tools and an outline of the proof
Section 7: proof of the main theorem
@@ Acknowledgments
We are grateful to Ian Agol, James Belk, Mark Bell, Joan Birman, Tara Brendle, Martin Bridson, Alex Eskin, Benson Farb, Juan González-Meneses, Toby Hall, Chris Leininger, Livio Liechti, Scott MacLachlan, Kasra Rafi, Saul Schleimer, Richard Webb, and Alex Wright for helpful conversations.
@@ 2.1. The ledger
The four train tracks $\tau_{1}$, $\tau_{2}$, $\tau_{3}$, and $\tau_{4}$ shown in the left-hand side of Figure 2 give a cell decomposition $Y$ of $\operatorname{MF}(\operatorname{D}_{3})$ as an integral cone complex. The top-dimensional cells are the $V_{i}=V(\tau_{i})$, and the faces are given by the rays through four curves shown. For further explanation of this picture, see the book by Farb and the first author [11, Chapter 15] (the train tracks there are slightly different). Here we have another difference from our general setup, as the train tracks here are not Dehn–Thurston train tracks; on the other hand these train tracks are especially suited to the braid group.
@@ 6.2. Proof of the fitting lemma
Before proving our main theorem, we require four technical lemmas, a criterion for carrying, an upper bound on the radius of a Dehn–Thurston cell structure, an upper bound for the radius of the set of vertex cycles in a Dehn–Thurston train track, and a linear algebraic lemma about projective sinks.
@@
Finally, we give a lemma that allows us to conclude the second statement of Theorem 1.1 from the first. As in the introduction, a vertex ray for a integral cone in an integral cone complex $Y$ is a vertex of the corresponding polyhedron in $\mathcal{P}(Y)$.

### v2

@@ 1.2. The algorithm
There are many examples of ledgers already in the literature. In his thesis [26], Penner gave a ledger in terms of Dehn–Thurston train track coordinates (explicit matrices are not given, but they can be derived from the computations there). Penner’s formulas contain minor mistakes, corrected in his subsequent paper [23]. There is still one other mistake, pointed out by Yandi Wu: in the second elementary move (Theorem A.2), the $\lambda$’s and $\kappa$’s should be switched [32].
@@
The complexity of the sink package algorithm is quadratic in the bit length of the entries of $A$. Indeed, the Zassenhaus algorithm converts the problem of computing a basis for an intersection into a row reduction problem, and the complexity of row reduction for integer matrices is known to be quadratic in the bit lengths of the entries; see, e.g. [33, Lecture X, Theorem 5].
@@
We are finally ready to state our corollary of Theorem 1.1, which gives the algorithm from the Main Theorem.
@@ 1.3. New tools and an outline of the proof
Section 7: proof of the main technical theorem
@@ Acknowledgments
We are grateful to Ian Agol, James Belk, Mark Bell, Joan Birman, Tara Brendle, Martin Bridson, Alex Eskin, Benson Farb, Juan González-Meneses, Toby Hall, Chris Leininger, Livio Liechti, Scott MacLachlan, Kasra Rafi, Roberta Shapiro, Saul Schleimer, Richard Webb, and Alex Wright for helpful conversations. We are also grateful to several anonymous referees for helpful comments and especially for pointing out the connection between our work and Thurston’s question.
@@ 2.1. The ledger
The four train tracks $\tau_{1}$, $\tau_{2}$, $\tau_{3}$, and $\tau_{4}$ shown in the left-hand side of Figure 4 give a cell decomposition $Y$ of $\operatorname{MF}(\operatorname{D}_{3})$ as an integral cone complex. The top-dimensional cells are the $V_{i}=V(\tau_{i})$, and the faces are given by the rays through four curves shown. For further explanation of this picture, see the book by Farb and the first author [11, Chapter 15] (the train tracks there are slightly different). Here we have another difference from our general setup, as the train tracks here are not Dehn–Thurston train tracks; on the other hand these train tracks are especially suited to the braid group. Also, this is a different structure on $\operatorname{MF}(\operatorname{D}_{3})\cong\operatorname{MF}(S_{0,4})$ given in the introduction; the calculation can be carried out just as well with that setup.
@@ 6.2. Proof of the fitting lemma
Before proving our main technical theorem, we require four technical lemmas, a criterion for carrying, an upper bound on the radius of a Dehn–Thurston cell structure, an upper bound for the radius of the set of vertex cycles in a Dehn–Thurston train track, and a linear algebraic lemma about projective sinks.
@@
Finally, we give a lemma that allows us to conclude the second statement of Theorem 1.1 from the first. As in the introduction, a vertex ray for an integral cone in an integral cone complex $Y$ is a vertex of the corresponding polyhedron in $\mathcal{P}(Y)$.